Find the interest rates earned on each of the following. Round each answer to two decimal places.

You borrow $700 and promise to pay back $763 at the end of 1 year.

%

You lend $700 and the borrower promises to pay you $763 at the end of 1 year.

%

You borrow $98,000 and promise to pay back $119,232 at the end of 5 years.

%

You borrow $9,000 and promise to make payments of $2,684.80 at the end of each year for 5 years.

%
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1.

763=700(1+r/100)^1

(763/700)=(1+r/100)

(1+r/100)=1.09

Hence r=(1.09-1)*100

=9%

2.

763=700(1+r/100)^1

(763/700)=(1+r/100)

(1+r/100)=1.09

Hence r=(1.09-1)*100

=9%

3.

119232=98000(1+r/100)^5

(119232/98000)^(1/5)=(1+r/100)

(1+r/100)=1.04

Hence r=(1.04-1)*100

=4%

4.

Let the rate be x%'

At this rate;present value of payments=$9000

9000=2684.8/1.0x+2684.8/1.0x^2+......+2684.8/1.0x^5

9000=2684.8[1/1.0x+1/1.0x^2+..........+1/1.0x^5]

[1/1.0x+1/1.0x^2+..........+1/1.0x^5]=(9000/2684.8)

=3.352(Approx)

Hence x=rate=15%(Looking at present value of annuity factor(15%,5 years)also].